# Deparametrization and Quantization of Scalar-Tensor Gravity and Its Cosmological Model

### v1

@@ III.3 Quantum Dynamics of the Scalar-Tensor Theory
$(\hat{C}^{*}(N)\cdot\Psi|\alpha,\phi\rangle=0,\ \ \ \ \forall\ \ |\alpha,\phi\rangle$ (III.57)
where we denote $|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle\equiv|\alpha,\phi\rangle$. By defining $\hat{H_{p}}(N):=-\hat{C_{2}^{\delta}}(N)-\hat{C_{3}^{\delta}}(N)$, Eq. (III.57) can be written as
$(\Psi|\hat{C_{1}}(N)|\alpha,\phi\rangle=(\Psi|\hat{H_{p}}(N)|\alpha,\phi\rangle$ (III.58)
Substituting Eq.(III.27) into Eq. (III.58), we have
$\frac{1}{2i\lambda_{0}}\!\Big(\!(\Psi|\alpha,\phi\!-\!\lambda_{0}N\rangle\!-\!(\Psi|\alpha,\phi\!+\!\lambda_{0}N\rangle\!\Big)\!=\!(\!\Psi|\hat{H_{p}}(N)|\alpha,\phi\rangle\!$ (III.59)
It should be noted that Eq. (III.59), which arises when performing polymer quantization with respect to $\pi$, can be regarded as the Schrödinger-like difference equation. It describes the discrete time evolution of physical states with respect to $\phi$, where $\hat{H_{p}}(N)$ plays the role of the physical Hamiltonian operator. Although it is challenging to provide a general solution of Eq. (III.59), it is possible to obtain specific solutions when applying this deparameterized framework of scalar-tenser theory to certain symmetry-reduced models. In the next section, we will apply this framework to a cosmological model and construct explicit dynamics based on the solution to the deparameterized Hamiltonian constraint equation.
@@
$\quad\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}^{\frac{2}{3}}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.60)
To quantize the deparametrized Hamiltonian constraint (III.7), we devide it into the following two terms:
$C_{1}^{\prime}(N)=\int d^{3}xN(x)\pi(x)$ (III.61)
@@
Note that the first term $C_{1}^{\prime}(N)$ has been quantized as (III.27). The second term $C_{2}^{\prime}(N)$ can be expressed as
$C_{2}^{\prime}(N)=\mp\lim_{\rho\rightarrow 0}\sum_{\square^{\rho}}\operatorname{sgn}\left(N_{\square^{\rho}}\right)\sqrt{T^{\prime}_{\square^{\rho}}}$ (III.62)
@@
$+\left(\int_{\square^{\rho}}d^{3}xN(x)\sqrt{\frac{\xi(\phi)}{\phi}}\sqrt{h}\right)^{2}$ (III.63)
Eq.(III.3) is written into six terms as $T^{\prime}_{\square^{\rho}}=\sum_{i=1}^{6}T^{\prime,i}_{\square^{\rho}}$. Notice that $T^{\prime,1}_{\square^{\rho}},T^{\prime,4}_{\square^{\rho}},T^{\prime,5}_{\square^{\rho}},T^{\prime,6}_{\square^{\rho}}$ take the same forms as the term $T^{1}_{\square^{\rho}},T^{3}_{\square^{\rho}},T^{4}_{\square^{\rho}},T^{5}_{\square^{\rho}}$ respectively in the sector of $\omega(\phi)\neq-\frac{3}{2}$, and hence have been quantized. To quantize the term $T_{\square^{\rho}}^{\prime,2}$, we use Eq. (III.31)–Eq. (III.34) to express it as
@@
$\!\times\!\left(\!\int\!d^{3}xN\!\delta_{v}^{k}\!\left\{\!A_{b}^{v},\left\{H^{E}(1),V\right\}\!\right\}\!\!\varepsilon_{lij}\!\varepsilon^{bef}\left\{A_{e}^{i},V\right\}\{A_{f}^{j},V\}\!\right)\!$ (III.64)
Then it can be regularized by using Eq. (III.37)–Eq. (III.38) as
@@
$\quad\times h_{s_{N}\left(\Delta^{\prime}\right)}\left\{h_{s_{N}\left(\Delta^{\prime}\right)}^{-1},V_{(v(\Delta^{\prime}),\epsilon)}\right\}\Big)$ (III.65)
@@
$\quad\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.66)
@@
$\!\hat{C}_{2}^{{}^{\prime}\delta}\!(\!N\!)\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!|\phi\rangle\!=\!\mp\!\sum_{v}\!sgn\!(\!N_{v}\!)\!\sqrt{\sum_{i=1}^{6}\widehat{T^{\prime,i,\delta}_{\square^{\rho,v}}}}\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!\cdot|\phi\rangle\!$ (III.68)
@@
$(\hat{C}^{{}^{\prime}*}(N)\cdot\Psi|\alpha,\phi\rangle=0,\ \ \ \ \forall\ \ |\alpha,\phi\rangle$ (III.69)
@@
$\frac{1}{2i\lambda_{0}}\!\Big(\!(\Psi|\alpha,\phi\!-\!\lambda_{0}N\rangle\!-\!(\Psi|\alpha,\phi\!+\!\lambda_{0}N\rangle\!\Big)\!=\!(\Psi|\hat{H_{p}^{{}^{\prime}}}(N)|\alpha,\phi\rangle$ (III.70)
@@ IV.2 Structure of the quantum theory
The deparametrized Hamiltonian constraint (III.3) can be expressed by the basics variables as

### v2

@@ III.3 Quantum Dynamics of the Scalar-Tensor Theory
$(\hat{C}^{*}(N)\cdot\Psi|\alpha,\phi\rangle=0,\ \ \ \ \forall\ \ |\alpha,\phi\rangle$ (III.58)
where we denote $|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle\equiv|\alpha,\phi\rangle$. By defining $\hat{H_{p}}(N):=-\hat{C_{2}^{\delta}}(N)-\hat{C_{3}^{\delta}}(N)$, Eq. (III.58) can be written as
$(\Psi|\hat{C_{1}}(N)|\alpha,\phi\rangle=(\Psi|\hat{H_{p}}(N)|\alpha,\phi\rangle$ (III.59)
Substituting Eq.(III.28) into Eq. (III.59), we have
$\frac{1}{2i\lambda_{0}}\!\Big(\!(\Psi|\alpha,\phi\!-\!\lambda_{0}N\rangle\!-\!(\Psi|\alpha,\phi\!+\!\lambda_{0}N\rangle\!\Big)\!=\!(\!\Psi|\hat{H_{p}}(N)|\alpha,\phi\rangle\!$ (III.60)
It should be noted that Eq. (III.60), which arises when performing polymer quantization with respect to $\pi$, can be regarded as the Schrödinger-like difference equation. It describes the discrete time evolution of physical states with respect to $\phi$, where $\hat{H_{p}}(N)$ plays the role of the physical Hamiltonian operator. Although it is challenging to provide a general solution of Eq. (III.60), it is possible to obtain specific solutions when applying this deparameterized framework of scalar-tenser theory to certain symmetry-reduced models. In the next section, we will apply this framework to a cosmological model and construct explicit dynamics based on the solution to the deparameterized Hamiltonian constraint equation.
@@
$\quad\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}^{\frac{2}{3}}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.61)
To quantize the deparametrized Hamiltonian constraint (III.8), we devide it into the following two terms:
$C_{1}^{\prime}(N)=\int d^{3}xN(x)\pi(x)$ (III.62)
@@
Note that the first term $C_{1}^{\prime}(N)$ has been quantized as (III.28). The second term $C_{2}^{\prime}(N)$ can be expressed as
$C_{2}^{\prime}(N)=\mp\lim_{\rho\rightarrow 0}\sum_{\square^{\rho}}\operatorname{sgn}\left(N_{\square^{\rho}}\right)\sqrt{T^{\prime}_{\square^{\rho}}}$ (III.63)
@@
$+\left(\int_{\square^{\rho}}d^{3}xN(x)\sqrt{\frac{\xi(\phi)}{\phi}}\sqrt{h}\right)^{2}$ (III.64)
Eq.(III.3) is written into six terms as $T^{\prime}_{\square^{\rho}}=\sum_{i=1}^{6}T^{\prime,i}_{\square^{\rho}}$. Notice that $T^{\prime,1}_{\square^{\rho}},T^{\prime,4}_{\square^{\rho}},T^{\prime,5}_{\square^{\rho}},T^{\prime,6}_{\square^{\rho}}$ take the same forms as the term $T^{1}_{\square^{\rho}},T^{3}_{\square^{\rho}},T^{4}_{\square^{\rho}},T^{5}_{\square^{\rho}}$ respectively in the sector of $\omega(\phi)\neq-\frac{3}{2}$, and hence have been quantized. To quantize the term $T_{\square^{\rho}}^{\prime,2}$, we use Eq. (III.32)–Eq. (III.35) to express it as
@@
$\!\times\!\left(\!\int\!d^{3}xN\!\delta_{v}^{k}\!\left\{\!A_{b}^{v},\left\{H^{E}(1),V\right\}\!\right\}\!\!\varepsilon_{lij}\!\varepsilon^{bef}\left\{A_{e}^{i},V\right\}\{A_{f}^{j},V\}\!\right)\!$ (III.65)
Then it can be regularized by using Eq. (III.38)–Eq. (III.39) as
@@
$\quad\times h_{s_{N}\left(\Delta^{\prime}\right)}\left\{h_{s_{N}\left(\Delta^{\prime}\right)}^{-1},V_{(v(\Delta^{\prime}),\epsilon)}\right\}\Big)$ (III.66)
@@
$\quad\cdot\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}\left[\hat{h}_{s_{N}\left(\Delta^{\prime}\right)}^{-1},\hat{V}_{(v(\Delta^{\prime}),\epsilon)}\right]\Big)\cdot|\beta,\vec{j},\vec{\iota}\rangle\otimes|\phi\rangle$ (III.68)
@@
$\!\hat{C}_{2}^{{}^{\prime}\delta}\!(\!N\!)\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!|\phi\rangle\!=\!\mp\!\sum_{v}\!sgn\!(\!N_{v}\!)\!\sqrt{\sum_{i=1}^{6}\widehat{T^{\prime,i,\delta}_{\square^{\rho,v}}}}\!\cdot\!|\beta,\!\vec{j},\!\vec{\iota}\rangle\!\otimes\!\cdot|\phi\rangle\!$ (III.69)
@@
$(\hat{C}^{{}^{\prime}*}(N)\cdot\Psi|\alpha,\phi\rangle=0,\ \ \ \ \forall\ \ |\alpha,\phi\rangle$ (III.70)
@@
$\frac{1}{2i\lambda_{0}}\!\Big(\!(\Psi|\alpha,\phi\!-\!\lambda_{0}N\rangle\!-\!(\Psi|\alpha,\phi\!+\!\lambda_{0}N\rangle\!\Big)\!=\!(\Psi|\hat{H_{p}^{{}^{\prime}}}(N)|\alpha,\phi\rangle$ (III.71)
@@ IV.2 Structure of the quantum theory
The deparametrized Hamiltonian constraint (III.4) can be expressed by the basics variables as